# Spins of Black Holes in X-ray Binaries and the Tension with the Gravitational Wave Measurements

### v2

@@ Abstract
We review current challenges in understanding the values and origin of the spins of black holes in binaries. Thanks to recent advances in astrophysical instrumentation, the spins can now be measured using both gravitational waves emitted by merging black holes and electromagnetic radiation from accreting X-ray binaries containing black holes. A key finding of the gravitational-wave observatories is that premerger black holes in binaries have low spin values, with an average dimensionless spin parameter of $a_{*}\sim$0.1–0.2, with 90% having $a_{*}\lesssim 0.6$. This implies that the natal spins of black holes are generally low, and the angular momentum transport in massive stars is efficient. On the other hand, most of the published spins in X-ray binaries are very high, $a_{*}\gtrsim 0.7$. In particular, this is the case for binaries with high-mass donors (potential progenitors of mergers), where their published spins range from 0.8 to 1.0. At the same time, their short lifetimes prevent significant spin-up by accretion. Those with low-mass donors could be spun-up to $a_{*}\gtrsim 0.5$ by conservative accretion. Spins $a_{*}\gtrsim 0.7$ can be achieved only if the donor initial masses were more than several solar masses, which remains unproven. However, the existing methods of spin measurements suffer from significant systematic errors. The method relying on relativistic X-ray line broadening is based on the separation of the observed spectra into incident and reflected ones, which is intrinsically highly uncertain. The method relying on spectral fitting of accretion disk continua uses models that predict the disk to be highly unstable, while stability is observed. Improved stable models yield disk temperatures higher than the standard models, and consequently predict lower spins. The published spin measurements in X-ray binaries are uncertain. The spins of the binaries with high-mass donors may be low, while those with low-mass donors have a broader spin distribution, ranging from low to high, including high spins as required to power relativistic jets.
@@ 2 Results from BH mergers
Mergers of BBHs can now be studied using GWs, which have allowed us to measure the spin of the premerger BHs. At least some BBHs (excluding those dynamically formed) are the final stages of binaries containing a BH and a high-mass donor. The latter includes currently accreting BHs, i.e., BH HMXBs. The results of those studies thus have important implications for comparing the populations of BBHs and BH XRBs and understanding their differences and similarities.
@@
One is that the distribution of the measured BH component masses in BBHs extends to much higher values, up to $\approx\!100{M}_{\odot}$ (The LIGO Scientific Collaboration et al., 2025a), than those observed in accreting BH XRBs. Until recently, the highest measured BH mass was that of Cyg X-1, $21.2\pm 2.2{M}_{\odot}$ (Miller-Jones et al., 2021). However, the recent investigation using detailed atmospheric (non-LTE) models of the donor by Ramachandran et al. (2025) found $M\approx(12.7$$17.8){M}_{\odot}$. We also note that a $33\pm 1{M}_{\odot}$ BH was recently discovered in a detached binary, Gaia BH3, within the Galaxy (Gaia Collaboration et al., 2024). However, as discussed in that work, the donor has a very low metallicity (allowing the BH progenitor to avoid strong mass loss in stellar winds), resulting in significantly higher expected BH masses in the halo compared to other Galactic components (Olejak et al., 2020). The system is associated with the metal-poor ED-2 stream, suggesting the role of dynamical interactions in its formation (Marín Pina et al., 2024). The mass range difference between EM and GW populations of BHs can be explained by both the observational selection effects, with mergers of heavier BHs giving stronger signals in GWs (Fishbach and Kalogera, 2022; Liotine et al., 2023), and the effect of the metallicity decreasing with the increasing redshift, which then reduces the rates of stellar winds, e.g., Belczyński et al. (2020).
@@
peaks at $a_{\rm eff}\approx 0.0$, with the inferred distribution skewed toward positive values extending only up to $\approx$0.5 (see figure 9 in The LIGO Scientific Collaboration et al. 2025b). Here, $M_{1,2}$ are the masses of the two BHs, and $\theta_{1,2}$ are the angles of the respective spin axes with respect to the binary orbital momentum axis. The leading spin-orbit coupling term determines the value of the effective spin in the post-Newtonian waveform fitted to the gravitational signal. Its values are measured with better precision than those of the individual BHs. The modelled distribution of the individual spins peaks within $a_{*}\approx 0.01$$0.23$ (at 90% credibility), and $\sim$90% of them are $a_{*}\lesssim$0.57 (The LIGO Scientific Collaboration et al., 2025b).
Figure 1: Distribution of the individual values of the lowest (denoted by $a_{A}$) and highest ($a_{B}$) spins of the components of premerger BH binaries. Adapted from The LIGO Scientific Collaboration et al. (2025b). See Section 2 for discussion.
Furthermore, the distributions of the higher and lower spins of the BHs in the premerger binaries differ significantly (as first noted by Biscoveanu et al. 2021). This is shown in Figure 1 (The LIGO Scientific Collaboration et al., 2025b), where we see that their distributions peak at $a_{A}=0.0$ and $a_{B}\approx 0.3$–0.4, respectively. A possible explanation for that is that the first-born BH rotates very slowly (at its low natal spin, see below), and then some process after its formation spins up the other BH. The scenario proposed by the recent literature involves tidal interactions of the first-born BH with a stripped progenitor of the other BH in a tight orbit, $\lesssim$1 day, which leads to the synchronous rotation of the BH progenitor (Qin et al., 2018; Mandel and Fragos, 2020; Ma and Fuller, 2023). These interactions occur after the formation of the first BH but before the formation of the second one (Olejak and Belczynski, 2021), as shown in the phase ‘Tidal spin-up’ of Figure 2.
@@
A promising scenario to make the less massive BH the slower spinning one involves mass ratio reversal during the first stable mass transfer, i.e., before the BH formation (Olejak and Belczynski, 2021; Broekgaarden et al., 2022; Adamcewicz et al., 2024; Olejak et al., 2024). Such a scenario is illustrated in Figure 2, adapted from Olejak et al. (2024). In it, Star 1 is heavier on the zero-age main sequence (ZAMS), and it efficiently transfers most of its mass (in its hydrogen envelope) to the companion Star 2 via Roche-lobe overflow (RLOF I). After that, it becomes a stripped helium-core star (Wolf-Rayet; WR) and the lighter of the two, while the binary becomes detached. Then, the WR star collapses to a BH with a low (natal) spin and with a much heavier companion. Next, the Star 2 evolves and starts to transfer mass to the BH companion (RLOF II). The accretion on it is, however, Eddington-limited. Therefore, it only slightly increases the BH mass during that short evolution stage, with most of the transferred mass of Star 2 being lost to outflows. Star 2 becomes then a WR star. This second stable mass transfer phase leads to a significant reduction of the binary separation due to a highly unequal mass ratio with the donor significantly more massive than the BH (see orbital evolution during mass transfer phase, e.g., in Tauris and van den Heuvel, 2023). The binary period is now so short that the tidal forces exerted by the BH on the companion spin it up, leading to its corotation with the binary. When Star 2 collapses to a BH, it has both the mass and the spin higher than the first-born BH. Finally, both BHs merge.
@@
The scenario with the fast spinning second-born BH via tidal spin-up has also been considered to facilitate the production of long-duration GRBs (see, e.g., Fryer et al. 2022; Bavera et al. 2022). The leading model for the production of those GRBs requires a collapse of a rapidly rotating massive star forming an accretion disk that, in turn, launches a strong jet powered by BH spin extraction (Blandford and Znajek, 1977). The propagating jet then produces the GRB (Woosley, 1993; MacFadyen and Woosley, 1999; MacFadyen et al., 2001).
Another possible mechanism that could spin up BHs is accretion. However, it would need to be several orders of magnitude above the Eddington limit in order to significantly affect the spin and masses of merging BHs (van Son et al., 2020). Such models are disfavorable on physical grounds, see the discussion of hypercritical accretion in the context of BH HMXBs in Section 4.2.1. Most currently used population models either adopt the Eddington limit or exceed it slightly. Still, the impact of hypothetical hypercritical accretion on the distribution of individual spins and properties of a population of GW sources has not yet been well explored.
@@
The probable cause of spin misalignment inferred for a fraction of GW sources is the contribution of systems formed through dynamical interactions, e.g., in dense stellar environments. Mergers originating from such environments are generally expected to exhibit an isotropic spin distribution, symmetric around an effective spin of $a_{\rm eff}\approx 0$ (Rodriguez et al., 2018; Arca Sedda et al., 2020). The spin distribution inferred from GW detections shows, however, an asymmetry, with a preference for positive effective spin values (Abbott et al., 2023; The LIGO Scientific Collaboration et al., 2025b); therefore, it cannot be fully explained by the dynamical channel itself. A scenario in which the population is dominated by isolated binaries, but includes a non-negligible fraction of dynamically formed systems, remains a plausible option. Interestingly, a distribution of effective spins consistent with current GW observations could also arise from mergers occurring in young massive clusters (Banerjee et al., 2023). In such environments, a significant fraction of binaries may preserve their initial spin alignment asymmetries, as their orientations are not entirely randomized by dynamical interactions.
@@
Recent studies by Tong et al. (2025) and Antonini et al. (2025) suggest the presence of a subpopulation of dynamically assembled second-generation mergers in the GW catalog, characterized by higher masses and larger effective spins (The LIGO Scientific Collaboration et al., 2025c). In particular, they argue that BBHs with component masses within the so-called pair-instability mass gap (Bond et al., 1984; Heger and Woosley, 2002; Fryer et al., 2001) involve a BH formed through a previous merger. They claim that such a scenario would place the lower edge of the gap at around 45–50 ${M}_{\odot}$. However, this interpretation faces some challenges, e.g., due to the existence of several events in which both components may reside within the mass gap (such as GW231028) and the observed prevalence of positive effective spins across the population, especially among the most massive systems (The LIGO Scientific Collaboration et al., 2025b, c). This distribution suggests preferential spin alignment in merging BBH binaries, contrary to expectations for the dynamical channel.
@@
One of the most intriguing events in GWTC-4 is GW231123 (The LIGO Scientific Collaboration et al., 2025a), which represents the most massive BBH merger detected to date. Its both component masses were around $100\,{M}_{\odot}$ each, and the BHs’ spins exhibit exceptionally large values of $0.90^{+0.10}_{-0.19}$ and $0.80^{+0.20}_{-0.51}$, standing out from the rest of the population. GW231123 poses significant challenges for classical dynamical hierarchical formation in star clusters. Such extremely high spins are unlikely for pure dynamically assembled systems (Fishbach et al., 2017; Gerosa and Berti, 2017), and the large recoil kicks threaten the retention of the merger remnant (Antonini and Rasio, 2016). Simulations indicate that only a small fraction of hierarchical mergers can reproduce such high spins since the progenitor BHs require well-aligned spins. This suggests that GW231123-like events may require stellar binary evolution either alone (Croon et al., 2025; Popa and de Mink, 2025; Gottlieb et al., 2025; Tanikawa et al., 2025), or with subsequent dynamical assembly (Stegmann et al., 2025). Alternatively, they might form in more exotic environments, such as active galactic nucleus disks (Delfavero et al., 2025; Bartos and Haiman, 2025; Kıroğlu et al., 2025).
@@
We note that the interpretations of GW observations are subject to significant biases and limitations due to, e.g., the restricted set of available templates (Divyajyoti et al., 2024). While the inferred fractions and the detailed shape of the spin parameters distribution are model-dependent, the conclusion that detected BHs have predominantly low spins ($a\sim$0.0–0.2) appears robust and has persisted across former observing runs (Abbott et al., 2023), including the most recent GWTC-4.0 (The LIGO Scientific Collaboration et al., 2025b), as well as among several alternative recent analyses, e.g., (Adamcewicz et al., 2025).
@@ 3 Results from X-ray binaries
All three aforementioned components are relativistically modified due to the Doppler and General Relativistic effects (Fabian et al., 1989; Dauser et al., 2013). The thermal disk spectra are also modified by radiative transfer in their top layers (Davis et al., 2005), and have the disk temperature increasing with $a_{*}$. Moreover, the value of the spin also affects the timing and polarization properties of the observed emission. Finally, the power of the jets launched by the accretion flow and the BH depends on the spin. We discuss these effects in detail below.
@@ 3.1.1 The reflection method
There are a few convolution codes available within the X-ray fitting suite xspec (Arnaud, 1996), in particular xilconv (Kolehmainen et al., 2011), which is based on the opacity tables of the rest-frame reflection code xilver (García et al., 2013) combined with the Green’s functions of Magdziarz and Zdziarski (1995). An alternative is rfxconv, which uses instead the opacity tables of the reflection code reflionx (Ross et al., 1999; Ross and Fabian, 2007). While the use of the convolution method with curved spectra is their major advantage, allowing us to estimate the reflection spectra from an arbitrary incident spectrum, their accuracy remains limited, see Ding et al. (2024). Furthermore, they are one-zone models, and the current version of xilconv uses the tables of the (older) version of xilver of García et al. (2013), in which the electron density is fixed at $10^{15}$ cm-3. The resulting rest-frame reflection spectra have to be then relativistically broadened, e.g., using the code relconv (Dauser et al., 2010).
@@
Another issue affecting the accuracy of the reflection method is the dependence of the reflection spectra on the irradiating flux, and through that, on the density of the reflector, since a fitting parameter of the reflection codes is the ionization parameter,
@@
where $F_{\rm{irr}}$ is the flux irradiating on the reflector and $n$ is the electron density. Thus, the higher the flux, the higher the density at a given fitted value of $\xi$. Many studies have been done assuming the density of $10^{15}$ cm-3 (García et al., 2013), which is applicable to AGN studies. The current version of xillver works for $n$ from $10^{15}$ to $10^{20}$ cm-3, but the disk densities in BH XRBs can be higher. There is an ongoing effort to further improve the accuracy of xillver (Ding et al., 2024). As Ding et al. (2024) note, a statistically excellent fit does not guarantee the correctness of the model parameters, in particular because there are many local minima in a multi-dimensional spectral fitting procedure. Another important caveat is that the reflecting slab in those codes is assumed to be heated only by the irradiating flux. However, disks in the soft state are strongly viscously heated, with that heating dominating the effect of the irradiation.
Another potentially important effect for measuring spins is the reflection of blackbody radiation returning to the disk owing to the BH gravity. This effect is included in some way in kerrbb and kerrbb2, which allow the user to switch on the self-irradiation, and it has also been included in recent reflection modelling. However, it is assumed that the disk is completely absorbing (i.e., with null albedo), which then only slightly increases the local temperatures. Schnittman and Krolik (2009) present calculations of the spectra and polarization of that component, assuming the disk is completely reflecting (i.e., with albedo =1=1). It then results in a weak and soft high-energy tail beyond the disk blackbody, with its amplitude and polarization degree increasing with the spin. The tail becomes noticeable only at $a_{*}\gtrsim 0.9$ (Schnittman and Krolik, 2009). Then, Taverna et al. (2021) and Mirzaev et al. (2024) presented calculations of that effect as a function of the disk ionization. On the other hand, Connors et al. (2020, 2021) included an irradiating blackbody component with free normalization and temperature in their reflection fits (i.e., without including their dependencies on the spin of Schnittman and Krolik 2009), and attributed it to the returning disk blackbody. Finally, Dauser et al. (2022) performed calculations of the returning coronal reflection.

### v3

@@ Abstract
We review current challenges in understanding the values and origins of the spins of black holes in binaries. Thanks to recent advances in astrophysical instrumentation, the spins can now be measured using both gravitational waves emitted by merging black holes and electromagnetic radiation from accreting X-ray binaries containing black holes. A key finding of the gravitational-wave observatories is that premerger black holes in binaries have low spin values, with an average dimensionless spin parameter of $a_{*}\sim$0.1–0.2, with 90% having $a_{*}\lesssim 0.6$. This implies that the natal spins of black holes are generally low, and the angular momentum transport in massive stars is efficient. On the other hand, most of the published spins in X-ray binaries are very high, $a_{*}\gtrsim 0.7$. In particular, this is the case for binaries with high-mass donors (potential progenitors of mergers), where their published spins range from 0.8 to 1.0. At the same time, their short lifetimes prevent significant spin-up by accretion. Those with low-mass donors could be spun-up to $a_{*}\gtrsim 0.5$ by conservative accretion. Spins $a_{*}\gtrsim 0.7$ can be achieved only if the donor initial masses were more than several solar masses, which remains unproven. However, the existing methods of spin measurements suffer from significant systematic errors. The method relying on relativistic X-ray line broadening is based on the separation of the observed spectra into incident and reflected ones, which is intrinsically highly uncertain. The method relying on spectral fitting of accretion disk continua uses models that predict the disk to be highly unstable, while stability is observed. Improved stable models yield disk temperatures higher than the standard models, and consequently predict lower spins. The published spin measurements in X-ray binaries are uncertain. The spins of the binaries with high-mass donors may be low, while those with low-mass donors have a broader spin distribution, ranging from low to high, including high spins as required to power relativistic jets.
@@ 2 Results from BH mergers
Mergers of BBHs can now be studied using GWs, which have allowed us to measure the spin of the premerger BHs. At least some BBHs (excluding those dynamically formed) are the final stages of binaries containing a BH and a high-mass donor. The latter includes currently accreting BHs, i.e., BH HMXBs. The results of these studies thus have important implications for comparing the populations of BBHs and BH XRBs, as well as for understanding their differences and similarities.
@@
One is that the distribution of the measured BH component masses in BBHs extends to much higher values, up to $\gtrsim\!100{M}_{\odot}$ (The LIGO Scientific Collaboration et al., 2025a), than those observed in accreting BH XRBs. Until recently, the highest measured BH mass in XRBs was that of Cyg X-1, $21.2\pm 2.2{M}_{\odot}$ (Miller-Jones et al., 2021). However, the recent investigation using detailed atmospheric (non-LTE) models of the donor by Ramachandran et al. (2025) found $M\approx(12.7$$17.8){M}_{\odot}$. We also note that a $33\pm 1{M}_{\odot}$ BH was recently discovered in a detached binary, Gaia BH3, within the Galaxy (Gaia Collaboration et al., 2024). However, as discussed in that work, the donor has a very low metallicity (allowing the BH progenitor to avoid strong mass loss in stellar winds), resulting in significantly higher expected BH masses in the halo compared to other Galactic components (Olejak et al., 2020). The system is associated with the metal-poor ED-2 stream, suggesting the role of dynamical interactions in its formation (Marín Pina et al., 2024). The mass range difference between EM and GW populations of BHs can be explained by both the observational selection effects, with mergers of heavier BHs giving stronger signals in GWs (Fishbach and Kalogera, 2022; Liotine et al., 2023), and the effect of the metallicity decreasing with the increasing redshift, which then reduces the rates of stellar winds, e.g., Belczyński et al. (2020).
@@
peaks at $a_{\rm eff}\approx 0.0$, with the inferred distribution skewed toward positive values extending only up to $\approx$0.5 (see figure 9 in The LIGO Scientific Collaboration et al. 2025b). Here, $M_{1,2}$ are the masses of the two BHs, and $\theta_{1,2}$ are the angles of the respective spin axes with respect to the binary orbital momentum axis. The leading spin-orbit coupling term determines the value of the effective spin in the post-Newtonian waveform fitted to the gravitational signal. Its values are measured with better precision than those of the individual BHs. The modeled distribution of the individual spins peaks within $a_{*}\approx 0.01$$0.23$ (at 90% credibility), and $\sim$90% of them are $a_{*}\lesssim$0.57 (The LIGO Scientific Collaboration et al., 2025b).
Figure 1: Distribution of the individual values of the lower (denoted by $a_{A}$) and higher ($a_{B}$) spins of the components of premerger BH binaries. Adapted from The LIGO Scientific Collaboration et al. (2025b). See Section 2 for discussion.
Furthermore, the distributions of the higher and lower spins of the BHs in the premerger binaries differ significantly (as first noted by Biscoveanu et al. 2021). This is shown in Figure 1 (The LIGO Scientific Collaboration et al., 2025b), where we see that their distributions peak at $a_{A}=0.0$ and $a_{B}\approx 0.3$–0.4, respectively. A possible explanation for that is that the first-born BH rotates very slowly (at its low natal spin, see below), and then some process after its formation spins up the other BH. The scenario proposed by recent literature involves tidal interactions of the first-born BH with a stripped progenitor of the other BH in a tight orbit, $\lesssim$1 day, which leads to the synchronous rotation of the BH progenitor (Qin et al., 2018; Mandel and Fragos, 2020; Ma and Fuller, 2023). These interactions occur after the formation of the first BH but before the formation of the second one (Olejak and Belczyński, 2021), as shown in the phase ‘Tidal spin-up’ of Figure 2.
@@
A promising scenario to make the less massive BH the slower spinning one involves mass ratio reversal during the first stable mass transfer, i.e., before the BH formation (Olejak and Belczyński, 2021; Broekgaarden et al., 2022; Adamcewicz et al., 2024; Olejak et al., 2024). Such a scenario is illustrated in Figure 2, adapted from Olejak et al. (2024). In it, Star 1 is heavier on the zero-age main sequence (ZAMS), and it efficiently transfers most of its mass (in its hydrogen envelope) to the companion Star 2 via Roche-lobe overflow (RLOF I). After that, it becomes a stripped helium-core star (Wolf-Rayet; WR) and the lighter of the two, while the binary becomes detached. Then, the WR star collapses to a BH with a low (natal) spin and with a much heavier companion. Next, the Star 2 evolves and starts to transfer mass to the BH companion (RLOF II). The accretion on it is, however, Eddington-limited. Therefore, it only slightly increases the BH mass during that short evolution stage, with most of the transferred mass of Star 2 being lost to outflows. Star 2 becomes then a WR star. This second stable mass transfer phase leads to a significant reduction of the binary separation due to a highly unequal mass ratio with the donor significantly more massive than the BH (see orbital evolution during mass transfer phase, e.g., in Tauris and van den Heuvel, 2023). The binary period is now so short that the tidal forces exerted by the BH on the companion spin it up, leading to its corotation with the binary. When Star 2 collapses to a BH, it has both the mass and the spin higher than the first-born BH. Finally, both BHs merge.
@@
The scenario with the fast-spinning second-born BH via tidal spin-up has also been considered to facilitate the production of long-duration GRBs (see, e.g., Fryer et al. 2022; Bavera et al. 2022). The leading model for the production of those GRBs requires a collapse of a rapidly rotating massive star forming an accretion disk that, in turn, launches a strong jet powered by BH spin extraction (Blandford and Znajek, 1977). The propagating jet then produces the GRB (Woosley, 1993; MacFadyen and Woosley, 1999; MacFadyen et al., 2001).
Another possible mechanism that could spin up BHs is accretion. However, it would need to be several orders of magnitude above the Eddington limit to significantly affect the spin and masses of merging BHs (van Son et al., 2020). Such models are disfavorable on physical grounds, see the discussion of hypercritical accretion in the context of BH HMXBs in Section 4.2.1. Most currently used population models either adopt the Eddington limit or exceed it slightly. Still, the impact of hypothetical hypercritical accretion on the distribution of individual spins and properties of a population of GW sources has not yet been well explored.
@@
The probable cause of spin misalignment inferred for a fraction of GW sources is the contribution of systems formed through dynamical interactions, e.g., in dense stellar environments. Mergers originating from such environments are generally expected to exhibit an isotropic spin distribution, symmetric around an effective spin of $a_{\rm eff}\approx 0$ (Rodriguez et al., 2018; Arca Sedda et al., 2020). The spin distribution inferred from GW detections shows, however, an asymmetry, with a preference for positive effective spin values (Abbott et al., 2023; The LIGO Scientific Collaboration et al., 2025b); therefore, it cannot be fully explained by the dynamical channel itself. A scenario in which the population is dominated by isolated binaries, but includes a non-negligible fraction of dynamically formed systems, remains a plausible option. Interestingly, a distribution of effective spins consistent with current GW observations could also arise from mergers occurring in young massive clusters (Banerjee et al., 2023). In such environments, a significant fraction of binaries may preserve their initial spin alignment asymmetries, since their orientations are not fully randomized by dynamical interactions.
@@
Recent studies by Tong et al. (2025) and Antonini et al. (2025) suggest the presence of a subpopulation of dynamically assembled second-generation mergers in the GW catalog, characterized by higher masses and larger effective spins (The LIGO Scientific Collaboration et al., 2025c). In particular, they argue that BBHs with component masses within the so-called pair-instability mass gap (Bond et al., 1984; Heger and Woosley, 2002; Fryer et al., 2001) involve a BH formed through a previous merger. They claim that such a scenario would place the lower edge of the gap at around 45–-$50{M}_{\odot}$. However, this interpretation faces some challenges, e.g., due to the existence of several events in which both components may reside within the mass gap (such as GW231028) and the observed prevalence of positive effective spins across the population, especially among the most massive systems (The LIGO Scientific Collaboration et al., 2025b, c). This distribution suggests preferential spin alignment in merging BBH binaries, contrary to expectations for the dynamical channel.
@@
One of the most intriguing events in GWTC-4 is GW231123 (The LIGO Scientific Collaboration et al., 2025a), which represents the most massive BBH merger detected to date. Its both component masses were around $100\,{M}_{\odot}$ each, and the BHs’ spins exhibit exceptionally large values of $0.90^{+0.10}_{-0.19}$ and $0.80^{+0.20}_{-0.51}$, standing out from the rest of the population. GW231123 poses significant challenges for classical dynamical hierarchical formation in star clusters. Such extremely high spins are unlikely for pure dynamically assembled systems (Fishbach et al., 2017; Gerosa and Berti, 2017), and the large recoil kicks threaten the retention of the merger remnant (Antonini and Rasio, 2016). Simulations indicate that only a small fraction of hierarchical mergers can reproduce such high spins since the progenitor BHs require well-aligned spins. This suggests that GW231123-like events may require stellar binary evolution either alone (Croon et al., 2025; Popa and de Mink, 2025; Gottlieb et al., 2025; Tanikawa et al., 2025), or with subsequent dynamical assembly (Stegmann et al., 2025). Alternatively, they might form in more exotic environments, such as active galactic nucleus (AGN) disks (Delfavero et al., 2025; Bartos and Haiman, 2025; Kıroğlu et al., 2025).
@@
We note that the interpretation of GW observations is subject to systematic biases and methodological limitations, including those arising from the restricted set of available waveform templates (Divyajyoti et al., 2024; Dhani et al., 2025). Although the inferred fractions and detailed shape of the underlying spin distribution depend on the chosen modeling framework, a robust feature across multiple observing runs is that detected BHs exhibit predominantly low spins ($a\sim 0.0$$0.2$) (Abbott et al., 2023). This conclusion persists in the most recent GWTC–4.0 results (The LIGO Scientific Collaboration et al., 2025b) and is consistent with several independent re-analyses of the current GW population (e.g., Adamcewicz et al., 2025).
Furthermore, Dhani et al. (2025) show that waveform inaccuracies can artificially inflate the inferred spin magnitudes, producing posteriors that favor high spins even when the true spins are modest. These biases are most pronounced for binaries with unequal masses and strong precession, implying that spin magnitudes may be systematically overestimated in precisely the systems where high spins would be most informative. Consequently, the already low incidence of high spins in the GW population may be even sparser once waveform-modeling systematics are fully accounted for.
@@ 3 Results from X-ray binaries
All three aforementioned components are relativistically modified due to the Doppler and General Relativistic effects (Fabian et al., 1989; Dauser et al., 2013). The thermal disk spectra are also modified by radiative transfer in their upper layers (Davis et al., 2005), and the disk temperature increases with $a_{*}$. Moreover, the spin value also affects the timing and polarization properties of the observed emission. Finally, the power of the jets launched by the accretion flow and the BH depends on the spin. We discuss these effects in detail below.
@@ 3.1.1 The reflection method
There are a few convolution codes available within the X-ray fitting suite xspec (Arnaud, 1996), in particular xilconv (Kolehmainen et al., 2011), which is based on the opacity tables of the rest-frame reflection code xilver (García et al., 2013) combined with the Green’s functions of Magdziarz and Zdziarski (1995). An alternative is rfxconv, which uses instead the opacity tables of the reflection code reflionx (Ross et al., 1999; Ross and Fabian, 2007). While the use of the convolution method with curved spectra is their major advantage, allowing us to estimate the reflection spectra from an arbitrary incident spectrum, their accuracy remains limited, see Ding et al. (2024). Furthermore, they are one-zone models, and the current version of xilconv uses the tables of the version of xilver of García et al. (2013) (before v.1.0) in which the electron density is fixed at $10^{15}$ cm-3. The resulting rest-frame reflection spectra have to be then relativistically broadened, e.g., using the code relconv (Dauser et al., 2010).
@@
Another issue affecting the accuracy of the reflection method is the dependence of the reflection spectra on the irradiating flux, and through that, on the density of the reflector, since a fitting parameter of the reflection codes is the ionization parameter
@@
where $F_{\rm{irr}}$ is the flux irradiating on the reflector and $n$ is the electron density. Thus, the higher the flux, the higher the density at a given fitted value of $\xi$. Many studies have been done assuming the density of $10^{15}$ cm-3 (García et al., 2013), which is applicable to AGN studies. The current version of xillver works for $n$ from $10^{15}$ to $10^{20}$ cm-3, but the disk densities in BH XRBs can be higher. There is an ongoing effort to improve the accuracy of xillver further (Ding et al., 2024). As Ding et al. (2024) note, a statistically excellent fit does not guarantee the correctness of the model parameters, in particular because there are many local minima in a multi-dimensional spectral fitting procedure. Another important caveat is that the reflecting slab in those codes is assumed to be heated only by the irradiating flux. However, disks in the soft state are strongly viscously heated, with that heating dominating the effect of the irradiation.
Another potentially important effect for measuring spins is the reflection of blackbody radiation returning to the disk owing to the BH gravity. This effect is included in some way in kerrbb and kerrbb2, which allow the user to switch on the self-irradiation, and it has also been included in recent reflection modeling. However, it is assumed that the disk is completely absorbing (i.e., with null albedo), which then only slightly increases the local temperatures. Schnittman and Krolik (2009) present calculations of the spectra and polarization of that component, assuming the disk is completely reflecting (i.e., with albedo =1=1). It then results in a weak and soft high-energy tail beyond the disk blackbody, with its amplitude and polarization degree increasing with the spin. The tail becomes noticeable only at $a_{*}\gtrsim 0.9$ (Schnittman and Krolik, 2009). Then, Taverna et al. (2021) presented calculations of that effect as a function of the disk ionization. On the other hand, Connors et al. (2020, 2021) included an irradiating blackbody component with free normalization and temperature in their reflection fits (i.e., without including their dependencies on the spin of Schnittman and Krolik 2009), and attributed it to the returning disk blackbody. Finally, Dauser et al. (2022) and Mirzaev et al. (2024) performed calculations of the returning coronal reflection.